Given the demand equation Q=250-50P . Determine the total revenue and marginal revenue functions. Does this situation occur in a perfectly competitive market? If yes why? If not why not?

1
Expert's answer
2022-02-14T15:00:29-0500

"Q=250-50P"

"T.R=Q\u00d7P"

"=(250\u00d750P)P"

"=250P-50P^{2}"

"MR=\frac{∆TR}{P}=250-100P"

Yes this is a perfectly competitive market. This is because the marginal revenue equation is an equation for a straight line.